At an absolute temperature of 300ka certain mass of gas occupies 1m3.find the volume of the gas when the absolute temperature is 375k
1
Expert's answer
2019-10-17T11:14:35-0400

The equation of states for an ideal gas says


"PV=nkT"

If the pressure is constant we obtain


"\frac{V_2}{V_1}=\frac{T_2}{T_1}"

So

"V_2=V_1\frac{T_2}{T_1}"

"V_2=1\:\rm m^3\times \frac{375\:\rm K}{300\:\rm K}=1.25\:\rm m^3"